# Filtering with Time-frequency Analysis: An Adaptive and Lightweight Model for Sequential Recommender Systems Based on Discrete Wavelet Transform

### v1

@@ 1 Introduction
Recommender Systems (RS) play a crucial role in our web experience Roy and Dutta (2022); Boka et al. (2024). Conventional RSs model user-item interactions in a static manner, which makes it challenging for them to model the dynamic user behaviors Roy and Dutta (2022); Wang et al. (2019). As users’ interests evolve dynamically over time, it is essential for Sequential Recommender Systems (SRS) to step in. SRSs recommend by modeling the dynamic user behavior patterns behind user-item interaction sequences. In contrast with traditional models, SRSs can capture the sequential patterns of user behavior, which makes it increasingly popular among researchers and platforms Chen et al. (2019); Lin et al. (2024).
@@
For the advantages of SRSs, recent years have witnessed a lot of SRS proposed based on different techniques Shin et al. (2024); Cui et al. (2024); Tan et al. (2016); Gao et al. (2023); Zhou et al. (2020). There are models based on Markov Chains He and McAuley (2016), Recurrent Neural Networks Tan et al. (2016) or Convolutional Neural Networks Tang and Wang (2018). With Transformer-based models showing their great performance in sequential tasks, many efforts have been put into Transformer-based SRSs Kang and McAuley (2018); Sun et al. (2019); Shin et al. (2024). However, more recent studies reveal that the item-to-item self-attention used in Transformer is a low-pass filter in frequency domain Park and Kim (2022); Raghu et al. (2021); Dong et al. (2021), which means it only cares about the global information but lacks consideration of local characteristics Xu et al. (2020). In the domain of SRS, the low-pass filtering nature of Transformer makes it tend to predict items according to the users’ global preferences, but in lack of consideration for users’ local interests and short-term abrupt interests Shin et al. (2024); Du et al. (2023). It is undemanding to capture users’ long-term interests, but it remains to be challenging to predict the users’ preferences by their short-term abrupt interests Shin et al. (2024). Fig.1 shows a case study about a user in Movielens-100K dataset whose interest switches from thriller films to children’s films and switches back later, which may be because of a family event or something and often occurs in other users. The sudden switches of user behavior cause abrupt changes in the sequence, from the perspective of frequency domain, the abrupt changes usually project into high frequency sub-signals, which are oversmoothed by self-attention due to its low-pass filtering nature Xu et al. (2020); Shin et al. (2024); Du et al. (2023). Further, some previous studies notice that the great deal of parameters brought by self-attention cause overfitting Zhou et al. (2022).
@@
To tackle the low-pass filtering nature and great deal of parameters of self-attention, we look for solutions in Digital Signal Processing (DSP). We notice that some previous studies have used the Fast Fourier Transform (FFT) to address the overfitting issue and the low-pass filtering characteristic of self-attention Zhou et al. (2022); Du et al. (2023); Shin et al. (2024). However, FFT analyzes the components of the signal from the whole range, and is primarily a global method for frequency extraction rather than a localized one. As we know, a person’s interests are neither kept forever nor hold same components all the time Huleihel et al. (2021). Traditional methods like FFT are uncompetitive in extract the components of the signal in a localized view, and are not appropriate potentially for extracting short-term abrupt interests of users. So we introduce Discrete Wavelet Transform Arfaoui et al. (2021); Mallat (1989) into the task of Sequential Recommendation (SR), which utilizes wavelet for localized frequency decomposition, corresponding to the need of localized interests extraction in the analysis of user-item interaction sequences.
@@
Wavelet Transform is a widely used technique in DSP Guo et al. (2022); Halidou et al. (2023). Since a user-item interaction sequence are discrete, we use the discrete version of Wavelet Transform. DWT decompose a signal into sub-signals within time windows of different locations and different scales, corresponding to different durations of user interests, which is an important motivation for our designation. By DWTRec, we model user interests with the analysis of components of user-item interaction sequence, avoiding complex item-to-item computing of sequence and effectively extract the high-frequency interests of users. To adaptively filter different types of interests of users, we design an adaptive time-frequency filter which learns the weights of different interests automatically from the datasets.
@@ 2.2 The Low-Pass Filtering Nature of Self-Attention
Recent studies reveal the self-attention is constantly a low-pass filter Park and Kim (2022); Raghu et al. (2021). In the domain of SR, the sudden switches of user interests cause abrupt changes to the signal trends, thus projecting into high-frequency information in frequency domain, which is usually smoothed out by self-attention Shin et al. (2024); Du et al. (2023); Xu et al. (2020).
@@ 2.3 Discrete Wavelet Transform
Discrete Wavelet Transform is the discrete version of Wavelet Transform, the main idea of DWT is to replace the infinite-length wave of Fourier Transform with a finite-length wave Arfaoui et al. (2021); Meyer (1992), so the technique can extract the sub-signals within a certain period. The designed finite-length wave is called wavelet. Along with each wavelet, there is Father Scaling Function and a Mother Wavelet Function, respectively for the extraction of overall value level of the signal and detailed signal characteristics in Wavelet Transform. In DWT the two functions are replaced by a group of discrete filters including $L,l$ corresponding to the Father Scaling Function and $H,h$ corresponding to Mother Wavelet Function. The computations of i-th decomposition results are as follows:
$\textit{A}^{\textit{i}}[\textit{n}]=\sum^{\textit{k}-1}_{\textit{k}=0}\textit{A}^{\textit{i}-1}[2\textit{n}-\textit{k}]\textit{L}[\textit{k}]$ (2)
$\textit{D}^{\textit{i}}[\textit{n}]=\sum^{\textit{k}-1}_{\textit{k}=0}\textit{A}^{\textit{i}-1}[2\textit{n}-\textit{k}]\textit{H}[\textit{k}]\$ (3)
For our implementation, we use Mallat algorithm Mallat (1989) to get better efficiency. A single level decomposition is defined as follows:
@@
Here, the * denotes linear convolution, $\downarrow$ means down-sampling which is a popular technique to reduce input length Zhou (2020). For multi-level decomposition (MWD), the process is defined as Algorithm 1. Notice that in the algorithm, the $A$ is discarded until the last loop, as a result of according to the theory of Wavelet Transform, only $A$ of the last single-level decomposition will be used for reconstruction. After running MWD, we get $A^{\gamma},D^{\gamma},D^{\gamma-1},...D^{1}$. From the perspective of DSP, the frequencies of the results grow larger from left to right. $A^{\gamma}$ records the value level of the input while $D$s focus on the detailed characteristics of the input. There are two key portraits of the filter group controlling this performance that $\sum_{i}L[i]^{2}=1$ and $\sum_{i}H[i]=0$. The different elements in $A,D$ mean frequency components within different time windows, the length of the time windows depend on the length of the filter and decomposition level. Fig. 2 visualizes the high and low pass decomposition filters of sym6, a popular wavelet in DSP.
@@ 2.4 Selection of Wavelet
An essential composition of DWT is the filter group offered by wavelets. A lot of efforts have been put into the selection of wavelets since the proposal of the method Silik et al. (2022); Jang et al. (2021); Guo et al. (2022). However, there is not a generalized rule that applies to different applications. We choose the widely used sym wavelets and coif wavelets for our implementation. For convenience and accuracy, we use pre-computed filters by Lee et al. (2019). The visualization of sym6 is shown in Fig. 2. As for the decomposition level, we recommend using a larger number when sequence get longer to reveal more details of the sequence.
@@ 2.5 Comparison between Fourier Transform and Wavelet Transform
Where $0\leq\textit{k}\leq\textit{N}-1$. The DFT aims to decompose a discrete sequence into different infinite waves with different frequencies. Contrastingly with DWT that decomposes inputs into combinations of finite waves with different frequencies. The DFT performs great when the components of the sequence are static, but in the domain of SR, where user interests are non-stationary Huleihel et al. (2021), the components of sequence are dynamic, the global analysis of DFT brings incompetent performance. Conversely, DWT gives consideration to the location of the sub-signals while telling the frequencies. As a result, DWT is our choice for building SRS.
@@ 3.1 Embedding Layer
In the first layer of DWTRec, we aim at transforming the different inputs into sequences a fixed length, and embedding them to get a unified and comprehensive representation of each sequence for the model. Define max sequence length as N. Firstly, we examine the length of input sequence $\textit{S}^{\textit{u}}$, if the length exceeds N, then only the rightmost N items will be retained, else if the length is shorter than N, we pad items with the ID of 0 on the left until the length is equal to N. We define a item embedding look-up matrix $\textit{M}\in\mathcal{R}^{|\textit{V}|\times\textit{d}}$ in this layer. The embedding vector of item v is $\textit{M}_{\textit{v}}$. The processed $\textit{S}^{\textit{u}}=\{\textit{v}^{\textit{u}}_{1},\textit{v}^{\textit{u}}_{2},...\textit{v}^{\textit{u}}_{N}\}$. We look up each $\textit{v}^{\textit{u}}_{\textit{i}}$ in the look-up matrix, and concatenate them to form $\textit{E}^{0}\in\mathcal{R}^{\textit{N}\times\textit{d}}$. To make the values of embedding sequence position-aware, we define a learnable position encoding matrix $\textit{P}\in\mathcal{R}^{\textit{N}\times\textit{d}}$ and add it to $\textit{E}^{0}$. To speed up the training process and regularize the model, two widely used techniques, layer normalization and dropout Ba (2016); Srivastava et al. (2014), are included in this layer. The computation is as follows:
@@ 3.2.1 Time-Frequency Filter Block
To prevent the gradient from vanishing when model gets deeper and get a more stable training process and better generalization ability, typical techniques i.e., skip connection, dropout and layer normalization are implemented in this layer He et al. (2016); Ba (2016); Srivastava et al. (2014):
@@ 4.1 Setup
To demonstrate the efficacy of DWTRec, we carefully select SR benchmark datasets from different domains with different average sequence lengths and sparsity. Detailed dataset characteristics are offered in Tab. 3. To make the results convincing, we select the representative SOTA baselines of 2022 Qiu et al. (2022), 2023 Du et al. (2023) and 2024 Shin et al. (2024) and directly take their official codes for model implementation.
@@ 4.1.1 Datasets
Datasets are as follows: Amazon Beauty, Sports, Toys (e-commerce), Yelp (business), LastFM (music), ML-1M (movie), we follow the preprocessing steps of Zhou et al. (2020, 2022); Shin et al. (2024).
(a) Average responds on different frequency levels.
Figure 6: Visualization of average spectrum of learned filters.
@@ 4.1.2 Baselines
  • CNN/RNN-based methods i.e., Caser Tang and Wang (2018) and GRU4Rec Tan et al. (2016).
  • Transformer-based methods including SASRec Kang and McAuley (2018) and BERT4Rec Sun et al. (2019).
  • FFT-based method including FMLPRec Zhou et al. (2022).
  • Transformer-based method with contrastive learning including DuoRec Qiu et al. (2022).
  • Transformer-based methods with FFT including two methods, FEARec Du et al. (2023) and BSARec Shin et al. (2024).
@@ 4.1.3 Metrics
We introduce widely used SR metrics Hit Rate (HR) @K and Normalized Discounted Cumulative Gain (NDCG) @K Busa-Fekete et al. (2012). The former focuses on the hit rate of the Top-K list, while the later focuses on the location of the ground truth item in the Top-K list.
@@ 4.2 Overall Performance Comparison
We follow Qiu et al. (2022); Du et al. (2023); Shin et al. (2024) on evaluating the methods over the whole item set for fair comparison Krichene and Rendle (2020). Detailed results are in Tab. 2.
@@ 4.5.1 Studies On Sequence Length
4.5.2 Studies On Same Model Setting
@@
Table 5: HR@5 results comparison under same model setting.
@@ 4.5.3 Model Complexity Analysis
In this section, we analyze the complexity of DWTRec. The complexity of parameters each layer is $O(\textit{N}\gamma\textit{d}+\textit{d})$. The complexity of extra space needed for runtime is $O(\textit{N}\gamma\textit{d}+\textit{N}\textit{d})$ per layer, where $\textit{N}\gamma\textit{d}$ comes from time-frequency filter and Nd comes from FFN. Contrastingly, the runtime time complexity of self-attention is $O(\textit{N}^{2}\textit{d})$, while the space complexity is $O(\textit{N}^{2})$.

### v5

@@ Abstract
Keywords:
Sequential Recommendation Recommender System Sparse Data Mining.
1 Introduction
@@
Recommender Systems (RS) play a crucial role in our web experience [25, 3]. Conventional RSs model user-item interactions in a static manner, which makes it challenging for them to model the dynamic user behaviors [25, 32]. As users’ interests evolve dynamically over time, it is essential for Sequential Recommender Systems (SRS) to step in. SRSs recommend by modeling the dynamic user behavior patterns behind user-item interaction sequences. In contrast with traditional models, SRSs can capture the sequential patterns of user behavior, which makes it increasingly popular among researchers and platforms [5, 19].
@@
For the advantages of SRSs, recent years have witnessed a lot of SRS proposed based on different techniques [26, 6, 30, 9, 35]. There are models based on Markov Chains [13], Recurrent Neural Networks [30] or Convolutional Neural Networks [31]. With Transformer-based models showing their great performance in sequential tasks, many efforts have been put into Transformer-based SRSs [16, 29, 26]. However, more recent studies reveal that the item-to-item self-attention used in Transformer is a low-pass filter in frequency domain [22, 24, 7], which means it only cares about the global information but lacks consideration of local characteristics [33]. In the domain of SRS, the low-pass filtering nature of Transformer makes it tend to predict items according to the users’ global preferences, but in lack of consideration for users’ local interests and short-term abrupt interests [26, 8]. It is undemanding to capture users’ long-term interests, but it remains to be challenging to predict the users’ preferences by their short-term abrupt interests [26]. Fig.1 shows a case study about a user in Movielens-100K dataset whose interest switches from thriller films to children’s films and switches back later, which may be because of a family event or something and often occurs in other users. The sudden switches of user behavior cause abrupt changes in the sequence, from the perspective of frequency domain, the abrupt changes usually project into high frequency sub-signals, which are oversmoothed by self-attention due to its low-pass filtering nature [33, 26, 8]. Further, some previous studies notice that the great deal of parameters brought by self-attention cause overfitting [36].
@@
To tackle the low-pass filtering nature and great deal of parameters of self-attention, we look for solutions in Digital Signal Processing (DSP). We notice that some previous studies have used the Fast Fourier Transform (FFT) to address the overfitting issue and the low-pass filtering characteristic of self-attention [36, 8, 26]. However, FFT analyzes the components of the signal from the whole range, and is primarily a global method for frequency extraction rather than a localized one. As we know, a person’s interests are neither kept forever nor hold same components all the time [14]. Traditional methods like FFT are uncompetitive in extract the components of the signal in a localized view, and are not appropriate potentially for extracting short-term abrupt interests of users. So we introduce Discrete Wavelet Transform [1, 20] into the task of Sequential Recommendation (SR), which utilizes wavelet for localized frequency decomposition, corresponding to the need of localized interests extraction in the analysis of user-item interaction sequences.
@@
Wavelet Transform is a widely used technique in DSP [10, 11]. Since a user-item interaction sequence are discrete, we use the discrete version of Wavelet Transform. DWT decompose a signal into sub-signals within time windows of different locations and different scales, corresponding to different durations of user interests, which is an important motivation for our designation. By DWTRec, we model user interests with the analysis of components of user-item interaction sequence, avoiding complex item-to-item computing of sequence and effectively extract the high-frequency interests of users. To adaptively filter different types of interests of users, we design an adaptive time-frequency filter which learns the weights of different interests automatically from the datasets.
@@ 2.2 The Low-Pass Filtering Nature of Self-Attention
Recent studies reveal the self-attention is constantly a low-pass filter [22, 24]. In the domain of SR, the sudden switches of user interests cause abrupt changes to the signal trends, thus projecting into high-frequency information in frequency domain, which is usually smoothed out by self-attention [26, 8, 33].
@@ 2.3 Discrete Wavelet Transform
Discrete Wavelet Transform is the discrete version of Wavelet Transform, the main idea of DWT is to replace the infinite-length wave of Fourier Transform with a finite-length wave [1, 21], so the technique can extract the sub-signals within a certain period. The designed finite-length wave is called wavelet. Along with each wavelet, there is Father Scaling Function and a Mother Wavelet Function, respectively for the extraction of overall value level of the signal and detailed signal characteristics in Wavelet Transform. In DWT the two functions are replaced by a group of discrete filters including $L,l$ corresponding to the Father Scaling Function and $H,h$ corresponding to Mother Wavelet Function. The computations of i-th decomposition results are as follows:
$\textit{A}^{\textit{i}}[\textit{n}]=\sum^{\textit{FL}-1}_{\textit{k}=0}\textit{A}^{\textit{i}-1}[2\textit{n}-\textit{k}]\textit{L}[\textit{k}]$ (2)
$\textit{D}^{\textit{i}}[\textit{n}]=\sum^{\textit{FL}-1}_{\textit{k}=0}\textit{A}^{\textit{i}-1}[2\textit{n}-\textit{k}]\textit{H}[\textit{k}]\$ (3)
For our implementation, we use Mallat algorithm [20] to get better efficiency. A single level decomposition is defined as follows:
@@
Here, the * denotes linear convolution, $\downarrow$ means down-sampling which is a popular technique to reduce input length [34]. For multi-level decomposition (MWD), the process is defined as Algorithm 1. Notice that in the algorithm, the $A$ is discarded until the last loop, as a result of according to the theory of Wavelet Transform, only $A$ of the last single-level decomposition will be used for reconstruction. After running MWD, we get $A^{\gamma},D^{\gamma},D^{\gamma-1},...D^{1}$. From the perspective of DSP, the frequencies of the results grow larger from left to right. $A^{\gamma}$ records the value level of the input while $D$s focus on the detailed characteristics of the input. There are two key portraits of the filter group controlling this performance that $\sum_{i}L[i]^{2}=1$ and $\sum_{i}H[i]=0$. The different elements in $A,D$ mean frequency components within different time windows, the length of the time windows depend on the length of the filter and decomposition level. Fig. 2 visualizes the high and low pass decomposition filters of sym6, a popular wavelet in DSP.
@@ 2.4 Selection of Wavelet
An essential composition of DWT is the filter group offered by wavelets. A lot of efforts have been put into the selection of wavelets since the proposal of the method [27, 15, 10]. However, there is not a generalized rule that applies to different applications. We choose the widely used sym wavelets and coif wavelets for our implementation. For convenience and accuracy, we use pre-computed filters by [18]. The visualization of sym6 is shown in Fig. 2. As for the decomposition level, we recommend using a larger number when sequence get longer to reveal more details of the sequence.
@@ 2.5 Comparison between Fourier Transform and Wavelet Transform
Where $0\leq\textit{k}\leq\textit{N}-1$. The DFT aims to decompose a discrete sequence into different infinite waves with different frequencies. Contrastingly with DWT that decomposes inputs into combinations of finite waves with different frequencies. The DFT performs great when the components of the sequence are static, but in the domain of SR, where user interests are non-stationary [14], the components of sequence are dynamic, the global analysis of DFT brings incompetent performance. Conversely, DWT gives consideration to the location of the sub-signals while telling the frequencies. As a result, DWT is our choice for building SRS.
@@ 3.1 Embedding Layer
In the first layer of DWTRec, we aim at transforming the different inputs into sequences a fixed length, and embedding them to get a unified and comprehensive representation of each sequence for the model. Define max sequence length as N. Firstly, we examine the length of input sequence $\textit{S}^{\textit{u}}$, if the length exceeds N, then only the rightmost N items will be retained, else if the length is shorter than N, we pad items with the ID of 0 on the left until the length is equal to N. We define a item embedding look-up matrix $\textit{M}\in\mathcal{R}^{|\textit{V}|\times\textit{d}}$ in this layer. The embedding vector of item v is $\textit{M}_{\textit{v}}$. The processed $\textit{S}^{\textit{u}}=\{\textit{v}^{\textit{u}}_{1},\textit{v}^{\textit{u}}_{2},...\textit{v}^{\textit{u}}_{N}\}$. We look up each $\textit{v}^{\textit{u}}_{\textit{i}}$ in the look-up matrix, and concatenate them to form $\textit{E}^{0}\in\mathcal{R}^{\textit{N}\times\textit{d}}$. To make the values of embedding sequence position-aware, we define a learnable position encoding matrix $\textit{P}\in\mathcal{R}^{\textit{N}\times\textit{d}}$ and add it to $\textit{E}^{0}$. To speed up the training process and regularize the model, two widely used techniques, layer normalization and dropout [2, 28], are included in this layer. The computation is as follows:
@@ 3.2.1 Time-Frequency Filter Block
To prevent the gradient from vanishing when model gets deeper and get a more stable training process and better generalization ability, typical techniques i.e., skip connection, dropout and layer normalization are implemented in this layer [12, 2, 28]:
@@ 4.1 Setup
To demonstrate the efficacy of DWTRec, we carefully select SR benchmark datasets from different domains with different average sequence lengths and sparsity. Detailed dataset characteristics are offered in Tab. 3. To make the results convincing, we select the representative SOTA baselines of 2022 [23], 2023 [8] and 2024 [26] and directly take their official codes for model implementation.
@@ 4.1.1 Datasets
Datasets are as follows: Amazon Beauty, Sports, Toys (e-commerce), Yelp (business), LastFM (music), ML-1M (movie), we follow the preprocessing steps of [35, 36, 26].
Figure 6: Visualization of average responses of learned filters of different levels.
@@ 4.1.2 Baselines
  • CNN/RNN-based methods i.e., Caser [31] and GRU4Rec [30].
  • Transformer-based methods including SASRec [16] and BERT4Rec [29].
  • FFT-based method including FMLPRec [36].
  • Transformer-based method with contrastive learning including DuoRec [23].
  • Transformer-based methods with FFT including two methods, FEARec [8] and BSARec [26].
@@ 4.1.3 Metrics
We introduce widely used SR metrics Hit Rate (HR) @K and Normalized Discounted Cumulative Gain (NDCG) @K [4]. The former focuses on the hit rate of the Top-K list, while the later focuses on the location of the ground truth item in the Top-K list.
@@ 4.2 Overall Performance Comparison
We follow [23, 8, 26] on evaluating the methods over the whole item set for fair comparison [17]. Detailed results are in Tab. 2.
@@ 4.5.1 Studies On Sequence Length
4.5.2 Studies On Same Setting of Hyper-parameters
@@
Table 5: HR@5 results comparison under same hyper-parameters.
@@ 4.5.3 Model Complexity Analysis
In this section, we analyze the complexity of DWTRec. The complexity of parameters each layer is $O(\textit{N}\gamma\textit{d}+\textit{d}^{2})$. The complexity of extra space needed for runtime is $O(\textit{N}\gamma\textit{d}+\textit{N}\textit{d})$ per layer, where $\textit{N}\gamma\textit{d}$ comes from time-frequency filter and Nd comes from FFN. Contrastingly, the runtime time complexity of self-attention is $O(\textit{N}^{2}\textit{d})$, while the space complexity is $O(\textit{N}^{2})$.